OpenStudy (anonymous):

Simplify. Express in simplest radical form. this not exactly a question but more like tell me if i am correct or not. (√3+√5)(√3+√5)=(√3)^2+(√5)^2=3+5=8?

OpenStudy (jdoe0001):

\(\bf (a+b)(a+b) \ne a^2 + b^2\\ \quad \\ (a+b)(a+b) = (a+b)^2\)

OpenStudy (anonymous):

=(√3+√5)^2?

OpenStudy (anonymous):

ohh which would make (3+5)^2 =64

OpenStudy (jdoe0001):

well... not quite.... \({\bf (\sqrt{3}+\sqrt{5})(\sqrt{3}+\sqrt{5})}\\ \quad \\ \begin{cases} \sqrt{3}\times\sqrt{3}+\\ \sqrt{3}\times \sqrt{5}+\\ \quad \\ \sqrt{5}\times \sqrt{3}+\\ \sqrt{5}\times \sqrt{5} \end{cases} \implies 3\quad +\quad \sqrt{15 }\quad +\quad\sqrt{15 }\quad +\quad 5 \)

OpenStudy (jdoe0001):

\(\bf \sqrt{3}+\sqrt{5} \ne 8\)

OpenStudy (anonymous):

 like that?